# FedLoGe: Joint Local and Generic Federated Learning under Long-tailed Data

### 17 Jan 2024 v1

@@ A.8 Projection Layer and Dot-regression Loss is Not Necessary for Fed-LoGe
A.9 Difference initialization methods of fixed classifier
@@ A.10 Discussion of Privacy
A.11 Comparison of Existing Methods
@@
1. Most existing methods in personalized federated learning (Collins et al. (2021); Li et al. (2021a); T Dinh et al. (2020); Zhang et al. (2022b); Dai et al. (2022)) utilize the global model to regularize or construct personalized models but do not evaluate them in a generic setup, failing to yield a strong global model without long-tail bias. This makes them unsuitable as a strong incentive for attracting new clients. In contrast, our approach learns models that excel in both personalized (GA-FR) and generic (FA-LR) setups without compromising either. 2. Existing methods in federated long-tailed learning (Shang et al. (2022b); Yang et al. (2023a); Shang et al. (2022a); Wang et al. (2022); Qian et al. (2023)), focusing only on calibrating long-tail bias for a single global model, overlook the aspect of representation learning (backbone) under federated long-tailed settings and fail to provide personalized models. For instance, CreFF involves post-classifier retraining but does not advance in boosting the backbone and does not offer personalized models. In contrast, Our FedLoGe achieves strong representation learning with SSE-C to support both global model and local models. 3. Existing work involving fixed classifiers has not delved deeply into the analysis of the initialization of these fixed classifiers. FedBABU (Oh et al. (2022)), for instance, merely employs a randomly initialized fixed linear layer. For FedETF (Li et al. (2023)), the Vallina fixed ETF classifier has suffered from feature degeneration, characterized by significant fluctuations between features and the class mean, while FedLoGo alleviates this issue by filtering out noisy smaller mean features (Fig. 1), making features more concise and effective (Fig. 3 (b,c) and Fig. 6). Besides, FedLoGe is less computationally expensive and easier to adapt to existing algorithms. DR loss is necessary for FedETF, which means some modifications are required when integrating with existing algorithms. Besides, FedETF needs a projection layer for optimization. For ResNet18 with four blocks, a 512x512 projection layer would be approximately 1MB, almost doubling the parameter size of Block 1, which is 0.56MB.

### 08 Mar 2024 v2

@@ A.9 Projection Layer and Dot-regression Loss is Not Necessary for Fed-LoGe
A.10 Difference initialization methods of fixed classifier
@@ A.11 Discussion of Privacy
A.12 Convergence Experiments
We have conducted experimental investigations into the convergence. Specifically, we explored the number of rounds required to reach an accuracy of 0.33 on CIFAR100 with an imbalance factor (IF) of 100 and $\alpha$ = 0.5, as documented in the table below. It is evident that the convergence rate of FedLoGe is almost on par with FedAvg and surpasses the current state-of-the-art (SOTA) methods.
Methods Convergence Round
FedETF 83
FedROD 68
FedLoGe (Ours) 62
Table 10: Comparison of Convergence Rounds for Different Methods.
A.13 Comparison of Existing Methods
@@
First, most existing methods in personalized federated learning (Collins et al. (2021); Li et al. (2021a); T Dinh et al. (2020); Zhang et al. (2022b); Dai et al. (2022)) utilize the global model to regularize or construct personalized models but do not evaluate them in a generic setup, failing to yield a strong global model without long-tail bias. This makes them unsuitable as a strong incentive for attracting new clients. In contrast, our approach learns models that excel in both personalized (GA-FR) and generic (FA-LR) setups without compromising either.
Second, existing methods in federated long-tailed learning (Shang et al. (2022b); Yang et al. (2023a); Shang et al. (2022a); Wang et al. (2022); Qian et al. (2023)), focusing only on calibrating long-tail bias for a single global model, overlook the aspect of representation learning (backbone) under federated long-tailed settings and fail to provide personalized models. For instance, CreFF involves post-classifier retraining but does not advance in boosting the backbone and does not offer personalized models. In contrast, Our FedLoGe achieves strong representation learning with SSE-C to support both global model and local models.
@@
Second, existing work involving fixed classifiers has not delved deeply into the analysis of the initialization of these fixed classifiers. FedBABU (Oh et al. (2022)), for instance, merely employs a randomly initialized fixed linear layer. For FedETF (Li et al. (2023)), the Vallina fixed ETF classifier has suffered from feature degeneration, characterized by significant fluctuations between features and the class mean, while FedLoGe alleviates this issue by filtering out noisy smaller mean features (Fig. 1), making features more concise and effective (Fig. 3 (a,b) and Fig. 6). Besides, FedLoGe is less computationally expensive and easier to adapt to existing algorithms. DR loss is necessary for FedETF, which means some modifications are required when integrating with existing algorithms. Besides, FedETF needs a projection layer for optimization. For ResNet18 with four blocks, a 512x512 projection layer would be approximately 1MB, almost doubling the parameter size of Block 1, which is 0.56MB.
A.14 Impact of Client Numbers and Participation Rates
To investigate the adaptability and efficiency of FedLoGe in varied federated learning environments, we designed experiments focusing on two main variables: the number of clients and their participation rates. The experiments spanned across CIFAR10/100-LT and ImageNet/iNat datasets, tailored to simulate environments ranging from full participation to more realistic, sporadic engagement.
For CIFAR10/100-LT, we utilized a setup involving 40 clients with 100% participation and 5 local epochs per training round. Conversely, ImageNet/iNat experiments were conducted with 20 clients at a 40% participation rate and 3 local epochs, introducing a scenario with reduced participation and increased data complexity.
Additionally, to explore the behavior of FedLoGe under more stringent conditions, we conducted experiments on CIFAR100-LT with an imbalance factor of 100 and a heterogeneity coefficient of 0.5, decreasing the participation rate to 30%. This setup aimed to test FedLoGe’s resilience and personalized performance against data imbalance and client heterogeneity. Extending our exploration to a cross-device FL setting, the client count was increased to 100 with a 30% participation rate, challenging FedLoGe with a highly distributed environment and testing its scalability and performance.
The experimental outcomes, detailed in Table 11, highlight FedLoGe’s superior performance across different settings, particularly showcasing its strengths in scenarios with varied client participation rates and numbers.
Method 40 clients (100% participate) 40 clients (30%) 100 clients (30%)
FedAvg GM0.3818/PM0.6214 0.3746/0.6205 0.3718/0.6384
FedRoD 0.3919/0.6919 0.3758/0.7156 0.3727/0.7665
FedETF 0.3825/0.6421 0.4180/0.7083 0.3847/0.7822
FedLoGe 0.4233/0.7285 0.4206/0.7524 0.4159/0.8008
Table 11: Impact of client diversity and participation rates on FedLoGe performance compared to other FL algorithms.
These results confirm FedLoGe’s robustness and adaptability, underlining its potential for broad applicability in federated learning scenarios characterized by varying degrees of client engagement and diversity.
A.15 Effectiveness of SSE-C
We opted for a learning rate of 0.0001 with 10,000 optimization steps on ImageNet-LT with 1000 classes. The order of magnitude for the norm variance is between $10^{-7}$ and $10^{-9}$. For the angles across each pair of classifier vectors, we visualized the cosine similarity in SSE-C in Fig. 7, indicating that our optimized SSE-C is almost identical to the dense ETF.
Figure 7: The heatmap of cosine similarity between classifier vectors in SSE-C.